# Takahiro Nishinakaa,b,c and Yutaka Yoshida d,e

### v1

@@ 1 Introduction
It has recently been pointed out that the $U(1)_{r}$ twisted compactification of 4d $\mathcal{N}=2$ SCFTs of Argyres–Douglas (AD) type [1, 2, 3, 4, 5] give rise to a series of 3d $\mathcal{N}=4$ superconformal field theories (SCFTs) that are realized as the IR fixed points of 3d $\mathcal{N}=2$ Chern–Simons (CS) matter theories [6, 7, 8, 9, 10]. In particular, the authors of [9] have proposed a general strategy to identify 3d theories corresponding to a given AD theory in four dimensions. A key ingredient in their discussions is the IR formula for the Schur index of the 4d AD theories [11], which is written in terms of the spectrum of 4d BPS states at generic points on the Coulomb branch. Since this BPS spectrum jumps when a wall of marginal stability is crossed, there are generally many different 3d $\mathcal{N}=2$ CS matter theories that are expected to flow to the same $\mathcal{N}=4$ fixed point in the infrared. Therefore, the wall-crossing phenomena for 4d BPS states lead to IR dualities for various 3d $\mathcal{N}=2$ CS matter theories [9]. When an AD theory has no flavor symmetry, the resulting 3d $\mathcal{N}=4$ theory at the fixed point is expected to be rank-zero, in the sense that both its Higgs and Coulomb branches are zero-dimensional. Supersymmetry enhancement of 3d $\mathcal{N}=2$ CS matter theories flowing to rank-zero theories in the infrared was first studied in [12] and further developed in [13, 14].
@@
In this paper, we apply the proposal of [9] to the $(A_{M-1},A_{N-1})$ theories for coprime $M$ and $N$, for which the corresponding VOAs in the sense of [15] are conjectured in [11] to be the vacuum sector of the $(M,M+N)$ $W_{M}$ minimal model. In particular, we identify a series of 3d $\mathcal{N}=2$ CS matter theories describing the $U(1)_{r}$ twisted $S^{1}$-reduction of $(A_{2},A_{N-1})$ and $(A_{3},A_{N-1})$ theories with $N$ coprime to three and four, respectively. To be more specific, we identify the gauge group, matter content, mixed CS-levels, the charges of matter fields, and the monopole superpotential that is expected to give rise to an IR supersymmetry enhancement to $\mathcal{N}=4$. We then confirm that the half index of these CS matter theories coincide with the Schur index of the parent 4d theories, $(A_{2},A_{N-1})$ and $(A_{3},A_{N-1})$. In addition, for $(A_{M-1},A_{N-1})$ with coprime $M\geq 5$ and $N$, we give a conjectural expression for the half-monodromy and the Schur index, from which one can read off the mixed CS-levels, the matter content and the charges of matter fields of the 3d $\mathcal{N}=2$ CS matter theory.
@@ 2.1 Half index from IR formula for 4d Schur index
where $(q)_{\infty}\coloneq\prod_{k=1}^{\infty}(1-q^{k})$, $r$ is the dimension of the Coulomb branch [^1] For the $(A_{M-1},A_{N-1})$ theory with coprime $M$ and $N$, r=(M1)(N1)/2r=(M-1)(N-1)/2., $\gamma$ runs over electro-magnetic charges of the stable BPS hypermultiplets in the chamber discussed above, and $X_{\gamma}$ is an element of the quantum torus algebra satisfying
@@
$E_{q}(X)\coloneq\prod_{k=0}^{\infty}(1+q^{k+\frac{1}{2}}X)^{-1}=\sum_{n=0}^{\infty}\frac{\big(-q^{\frac{1}{2}}X\big)^{n}}{(q)_{n}}~,$ (2.4)
where $(q)_{n}\equiv\prod_{k=1}^{n}(1-q^{k})$. Note that $E_{q}(X_{\gamma_{1}})$ and $E_{q}(X_{\gamma_{2}})$ do not commute unless $\langle\gamma_{1},\gamma_{2}\rangle=0$, and therefore the order of $E_{q}(X_{\gamma})$ must be specified in (2.2); the order is taken according to the value of $\text{arg}\,Z(\gamma)$ where $Z(\gamma)$ is the central charge of the 4d $\mathcal{N}=2$ super-Poincaré algebra. The quantity $\overline{S(q)}$ is obtained from $S(q)$ by replacing all the electro-magnetic charges $\gamma$ to their charge-conjugates $\gamma$. The trace in (2.2) is then defined by [^2] When the theory has a flavor symmetry, Eq. (2.5) is slightly modified.
@@ 2.2 Ellipsoid partition function from IR formula
up to an overall constant. Here, $\Phi_{b}$ is the Faddeev quantum dilogarithm [30, 31, 32] (we follow the notation of [32] here):
@@
Note that the factor $\Phi_{\sf b}(\frac{i({\sf b}+{\sf b}^{-1})}{2}-p)$ in (2.10) is almost identical to the contribution from a chiral multiplet to the localization formula for the ellipsoid partition function, where the chiral multiplet is assigned with R-charge zero and is coupled to a $U(1)$ vector multiplet whose real scalar is $p$. This implies that (2.6) is identified as the ellipsoid partition function of an $\mathcal{N}=2$ CS matter theory, where $e^{-\pi ip^{2}}$ in (2.10) is a part of the CS level of a $U(1)$ gauge group corresponding to $p$. The precise identification given in the next sub-section of (2.6) with the ellipsoid partition function leads to the the same CS matter theory obtained by the Schur index Specifically, the we will that the same linear combination of topological $U(1)$ symmetries can be read off from the Fayet-Iliopoulos (FI) term.
@@ 2.3 General formula for 3d CS matter theory from IR formulas
$\mathcal{I}^{\text{4d}}$ $=(q)^{L-N}_{\infty}\sum_{n_{1},\cdots,n_{N}=0}^{\infty}\;\sum_{(n\cdot Q)_{1}\geq 0,\cdots,(n\cdot Q)_{L}\geq 0}\!\!\frac{q^{\frac{1}{2}\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}}(-q^{\frac{1}{2}})^{\sum_{a=1}^{N}\sum_{i=1}^{L}Q_{a,i}n_{a}}}{\prod_{i=1}^{L}(q)_{\sum_{a=1}^{N}n_{a}Q_{a,i}}}\,$, (2.14)
@@
$L:=2L^{\prime}\,,\quad N:=2L^{\prime}-2r\,.$ (2.16)
@@
where $(a)_{\infty}:=(a;q)_{\infty}=\prod_{k=0}^{\infty}(1-aq^{k})$, and $s_{\sf b}$ is the double sine function given by (A.8). By comparing these expressions with the localization formulas (A.1) and (A.5), we find that the Schur index and $\mathcal{S}^{\rm 3d}_{\sf b}$ agree with the half index and the ellipsoid partition function of a 3d $\mathcal{N}=2$ CS matter theory with the $U(1)^{N}$ gauge group coupled to $L$ chiral multiplets of vanishing R-charge, with gauge charges $Q_{ai}$ for $a=1,\ldots,N$ and $i=1,\ldots,L$. The effective gauge CS level is $K_{ab}$.
In particular, if the fugacities $x_{a}$ for the topological $U(1)$ symmetries and the Fayet–Iliopoulos (FI) parameters $\xi_{a}$ are turned off, i.e. $x_{a}=1$ and $\xi_{a}=0$, we find that the gauge–R-symmetry mixed CS levels is
@@ 2.4.2 $(A_{1},A_{2N})$ theories
By replacing $E_{q}(X_{\gamma})$ in (2.39) with $\Phi(x_{\gamma})$, one can identify the ellipsoid partition function (2.6) of the twisted $S^{1}$-compactification of the $(A_{1},A_{2N})$ theory. The result is expressed as
@@ 4.1 Schur index for $(A_{2},A_{N-1})$
$\mathcal{I}^{\text{4d}}$ $=(q;q)_{\infty}^{2(N-1)}\sum_{n^{(1)}_{i},n^{(2)}_{i}=0}^{\infty}\sum_{k_{i},\widetilde{k}_{i}=0}^{\text{min}(n^{(1)}_{i},n^{(2)}_{i})}q^{\sum_{i=1}^{N-2}(n^{(1)}_{i}n^{(1)}_{i+1}+n^{(2)}_{i}n^{(2)}_{i+1}-k_{i}k_{i+1}-\widetilde{k}_{i}\widetilde{k}_{i+1})-\sum_{i=1}^{N-1}(n^{(1)}_{i}n^{(2)}_{i}-\frac{1}{2}k_{i}^{2}-\frac{1}{2}\widetilde{k}_{i}^{2})}$
$\qquad\times\frac{\prod_{i=1}^{N-1}q^{n^{(1)}_{i}}q^{n^{(2)}_{i}}(-q^{-\frac{1}{2}})^{k_{i}}(-q^{-\frac{1}{2}})^{\widetilde{k}_{i}}}{(q;q)_{n^{(1)}_{i}-k_{i}}(q)_{n^{(2)}_{i}-k_{i}}(q)_{k_{i}}(q)_{n^{(1)}_{i}-\widetilde{k}_{i}}(q)_{n^{(2)}_{i}-\widetilde{k}_{i}}(q)_{\widetilde{k}_{i}}}~.$ (4.4)
@@ 4.2 Half index and CS matter theory
and label the chiral multiplets by $\phi^{(1)}_{i},\phi^{(2)}_{i},\widetilde{\phi}^{(1)}_{i},\widetilde{\phi}^{(2)}_{i},\varphi_{i},\widetilde{\varphi}_{i}$ (for $i=1,\ldots,N-1$), as depicted in Figure 7. We assume that the classical R-charges of all these chiral multiplets are zero.
@@
$\widetilde{K}_{ab}:=\begin{cases}1&\text{if }a=b-1,\quad b=1,\ldots,2N-1,\\[2.0pt] -1&\text{if }a=b-1,\quad b=2N-1,\ldots,4(N-1),\\[2.0pt] -1&\text{if }a=b-N+1,\quad b=1,\ldots,N-1,\\[2.0pt] \dfrac{1}{2}&\text{if }a=b,\quad b=2N-1,\ldots,4(N-1),\\[2.0pt] 0&\text{otherwise}.\end{cases}$ (4.7)
@@
Figure 7: The quiver diagram for the $i$-th set of gauge and matter multiplets included in the CS-matter theory for $(A_{2},A_{N-1})$ (for NN coprime to 33). Each circle stands for a $U(1)$ gauge group. Each arrow connecting two circles stands for a bifundamental chiral multiplets, and each arrow from a box to a circle stands for a chiral multiplet charged under the $U(1)$ corresponding to the circle. The subscript of $U(1)$ in each circle stands for the fugacity for the corresponding topological symmetry. This diagram can be read off from (4.4).
@@ 4.3 Monopole superpotential
Table 2: The magnetic charge for monopole operators in the $i$-th sector, $i=1,\dots,N-1$. The monopole operators in the $i$-th sector carry no charge under the gauge groups of the other sectors..
The non-vanishing fugacities (4.9) for the topological symmetries imply that the correct R-symmetry is a linear combination of the classical R-symmetry and the topological symmetries. This suggests that a superpotential involving monopole operators is turned on. From the formula for the gauge charge of monopole operator given in appendix C, we find that the following six types of dressed monopole operators satisfy the gauge-invariance condition $Q_{a}[\mathcal{O}_{(n,m)}]=0$:
@@
Note that, when the above terms are added to the superpotential, the classical R-symmetry is broken since the monopole operators are charged under the R-symmetry (C.2). However, the following linear combination of the UV R-symmetry and topological symmetries is preserved:
@@
$W$ $:=\sum_{i=1}^{N-1}\Bigl({\phi}^{(1)}_{i-1}{V}^{(1)}_{i}\phi^{(1)}_{i+1}+{\phi}^{(2)}_{i-1}{V}^{(2)}_{i}\phi^{(2)}_{i+1}+\widetilde{\phi}^{(1)}_{i-1}\widetilde{V}_{i}^{(1)}\widetilde{\phi}^{(1)}_{i+1}$
@@ 4.4 Residual global symmetry
$W$ $\to\sum_{k=1}^{N-1}\left(e^{i\beta_{k}}\varphi_{k-1}V_{k}\,\varphi_{k+1}+e^{i\left(-\frac{\zeta^{(1)}_{k}+\zeta^{(2)}_{k}}{2}+\widetilde{\beta}_{k}\right)}\widetilde{\varphi}_{k-1}\widetilde{V}_{k}\,\widetilde{\varphi}_{k+1}\right)~,$ (4.22)
$+\sum_{k=1}^{N-1}\left(e^{i\left(-\frac{\zeta_{k}^{(1)}}{2}-\alpha^{(1)}_{k}-\beta_{k}\right)}\phi_{k-1}^{(1)}V_{k}^{(1)}\phi_{k+1}^{(1)}+e^{i\left(-\frac{\zeta^{(2)}_{k}}{2}-\alpha^{(2)}_{k}-\beta_{k}\right)}\phi_{k-1}^{(2)}V_{k}^{(2)}\phi_{k+1}^{(2)}\right)~,$ (4.23)

### v2

@@ 1 Introduction
It has recently been pointed out that the $U(1)_{r}$ twisted compactification of 4d $\mathcal{N}=2$ superconformal field theories (SCFTs) of Argyres–Douglas (AD) type [1, 2, 3, 4, 5] give rise to a series of 3d $\mathcal{N}=4$ SCFTs that are realized as the IR fixed points of 3d $\mathcal{N}=2$ Chern–Simons (CS) matter theories [6, 7, 8, 9, 10]. In particular, the authors of [9] have proposed a general strategy to identify 3d theories corresponding to a given AD theory in four dimensions. A key ingredient in their discussions is the IR formula for the Schur index of the 4d AD theories [11], which is written in terms of the spectrum of 4d BPS states at generic points on the Coulomb branch. Since this BPS spectrum jumps when a wall of marginal stability is crossed, there are generally many different 3d $\mathcal{N}=2$ CS matter theories that are expected to flow to the same $\mathcal{N}=4$ fixed point in the infrared. Therefore, the wall-crossing phenomena for 4d BPS states lead to IR dualities for various 3d $\mathcal{N}=2$ CS matter theories [9]. When an AD theory has no flavor symmetry, the resulting 3d $\mathcal{N}=4$ theory at the fixed point is expected to be rank-zero, in the sense that both its Higgs and Coulomb branches are zero-dimensional. Supersymmetry enhancement of 3d $\mathcal{N}=2$ CS matter theories flowing to rank-zero theories in the infrared was first studied in [12] and further developed in [13, 14].
@@
In this paper, we apply the proposal of [9] to the $(A_{M-1},A_{N-1})$ theories for coprime $M$ and $N$, for which the corresponding VOAs in the sense of [15] are conjectured in [11] to be the vacuum sector of the $(M,M+N)$ $W_{M}$ minimal model. In particular, we identify a series of 3d $\mathcal{N}=2$ CS matter theories describing the $U(1)_{r}$ twisted $S^{1}$-reduction of $(A_{2},A_{N-1})$ and $(A_{3},A_{N-1})$ theories with $N$ coprime to three and four, respectively. To be more specific, we identify the gauge group, matter content, mixed CS levels, the charges of matter fields, and the monopole superpotential that is expected to give rise to an IR supersymmetry enhancement to $\mathcal{N}=4$. We then confirm that the half index of these CS matter theories coincide with the Schur index of the parent 4d theories, $(A_{2},A_{N-1})$ and $(A_{3},A_{N-1})$. In addition, for $(A_{M-1},A_{N-1})$ with coprime $M\geq 5$ and $N$, we give a conjectural expression for the half-monodromy and the Schur index, from which one can read off the mixed CS levels, the matter content and the charges of matter fields of the 3d $\mathcal{N}=2$ CS matter theory.
@@ 2.1 Half index from IR formula for 4d Schur index
where $(q)_{\infty}\equiv(q;q)_{\infty}=\prod_{k=1}^{\infty}(1-q^{k})$, $r$ is the dimension of the Coulomb branch [^1] For the $(A_{M-1},A_{N-1})$ theory with coprime $M$ and $N$, r=(M1)(N1)/2r=(M-1)(N-1)/2., $\gamma$ runs over electro-magnetic charges of the stable BPS hypermultiplets in the chamber discussed above, and $X_{\gamma}$ is an element of the quantum torus algebra satisfying
@@
$E_{q}(X)\equiv\prod_{k=0}^{\infty}(1+q^{k+\frac{1}{2}}X)^{-1}=\sum_{n=0}^{\infty}\frac{\big(-q^{\frac{1}{2}}X\big)^{n}}{(q)_{n}}~,$ (2.4)
where $(q)_{n}\equiv(q;q)_{n}=\prod_{k=1}^{n}(1-q^{k})$. Note that $E_{q}(X_{\gamma_{1}})$ and $E_{q}(X_{\gamma_{2}})$ do not commute unless $\langle\gamma_{1},\gamma_{2}\rangle=0$, and therefore the order of $E_{q}(X_{\gamma})$ must be specified in (2.2); the order is taken according to the value of $\text{arg}\,Z(\gamma)$ where $Z(\gamma)$ is the central charge of the 4d $\mathcal{N}=2$ super-Poincaré algebra. The quantity $\overline{S(q)}$ is obtained from $S(q)$ by replacing all the electro-magnetic charges $\gamma$ to their charge-conjugates $\gamma$. The trace in (2.2) is then defined by [^2] When the theory has a flavor symmetry, Eq. (2.5) is slightly modified.
@@ 2.2 Ellipsoid partition function from IR formula
up to an overall constant. Here, $\Phi_{\sf b}$ is the Faddeev quantum dilogarithm [30, 31, 32] (we follow the notation of [32] here):
@@
Note that the factor $\Phi_{\sf b}(\frac{i({\sf b}+{\sf b}^{-1})}{2}-p)$ in (2.10) is almost identical to the contribution from a chiral multiplet to the localization formula for the ellipsoid partition function, where the chiral multiplet is assigned with R-charge zero and is coupled to a $U(1)$ vector multiplet whose real scalar is $p$. This implies that (2.6) is identified as the ellipsoid partition function of an $\mathcal{N}=2$ CS matter theory, where $e^{-\pi ip^{2}}$ in (2.10) is a part of the CS level of a $U(1)$ gauge group corresponding to $p$. The precise identification given in the next sub-section of (2.6) with the ellipsoid partition function leads to the the same CS matter theory obtained by the Schur index. Specifically, we will see that one can read off from the half-index and the ellipsoid partition function that the same linear combination of the reference R-charge and topological $U(1)$ charges is allowed as a symmetry.
@@ 2.3 General formula for 3d CS matter theory from IR formulas
$\mathcal{I}^{\text{4d}}$ $=(q)^{L-N}_{\infty}\sum_{n_{1},\cdots,n_{N}\in\mathbb{Z}\atop(n\cdot Q)_{1}\geq 0,\cdots,(n\cdot Q)_{L}\geq 0}\!\!\frac{q^{\frac{1}{2}\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}}(-q^{\frac{1}{2}})^{\sum_{a=1}^{N}\sum_{i=1}^{L}Q_{a,i}n_{a}}}{\prod_{i=1}^{L}(q)_{\sum_{a=1}^{N}n_{a}Q_{a,i}}}\,$, (2.14)
@@
$L\equiv 2L^{\prime}\,,\quad N\equiv 2L^{\prime}-2r\,.$ (2.16)
@@
where $(a)_{\infty}\equiv(a;q)_{\infty}=\prod_{k=0}^{\infty}(1-aq^{k})$, and $s_{\sf b}$ is the double sine function given by (A.8). By comparing these expressions with the localization formulas (A.1) and (A.5), we find that the Schur index and $\mathcal{S}^{\rm 3d}_{\sf b}$ agree with the half index and the ellipsoid partition function of a 3d $\mathcal{N}=2$ CS matter theory with the $U(1)^{N}$ gauge group coupled to $L$ chiral multiplets of vanishing R-charge, with gauge charges $Q_{a,i}$ for $a=1,\ldots,N$ and $i=1,\ldots,L$. The effective gauge CS level is $K_{ab}$.
In particular, if the fugacities $x_{a}$ for the topological $U(1)$ symmetries and the Fayet–Iliopoulos (FI) parameters $\xi_{a}$ are turned off, i.e. $x_{a}=1$ and $\xi_{a}=0$, we find that the gauge–R-symmetry mixed CS level is
@@ 2.4.2 $(A_{1},A_{2N})$ theories
By replacing $E_{q}(X_{\gamma})$ in (2.39) with $\Phi_{\sf b}(x_{\gamma})$, one can identify the ellipsoid partition function (2.6) of the twisted $S^{1}$-compactification of the $(A_{1},A_{2N})$ theory. The result is expressed as
@@ 4.1 Schur index for $(A_{2},A_{N-1})$
$\mathcal{I}^{\text{4d}}$ $=(q)_{\infty}^{2(N-1)}\sum_{n^{(1)}_{i},n^{(2)}_{i}=0}^{\infty}\sum_{k_{i},\widetilde{k}_{i}=0}^{\text{min}(n^{(1)}_{i},n^{(2)}_{i})}q^{\sum_{i=1}^{N-2}(n^{(1)}_{i}n^{(1)}_{i+1}+n^{(2)}_{i}n^{(2)}_{i+1}-k_{i}k_{i+1}-\widetilde{k}_{i}\widetilde{k}_{i+1})-\sum_{i=1}^{N-1}(n^{(1)}_{i}n^{(2)}_{i}-\frac{1}{2}k_{i}^{2}-\frac{1}{2}\widetilde{k}_{i}^{2})}$
$\qquad\times\frac{\prod_{i=1}^{N-1}q^{n^{(1)}_{i}}q^{n^{(2)}_{i}}(-q^{-\frac{1}{2}})^{k_{i}}(-q^{-\frac{1}{2}})^{\widetilde{k}_{i}}}{(q)_{n^{(1)}_{i}-k_{i}}(q)_{n^{(2)}_{i}-k_{i}}(q)_{k_{i}}(q)_{n^{(1)}_{i}-\widetilde{k}_{i}}(q)_{n^{(2)}_{i}-\widetilde{k}_{i}}(q)_{\widetilde{k}_{i}}}~.$ (4.4)
@@ 4.2 Half index and CS matter theory
and label the chiral multiplets by $\phi^{(1)}_{i},\phi^{(2)}_{i},\widetilde{\phi}^{(1)}_{i},\widetilde{\phi}^{(2)}_{i},\varphi_{i},\widetilde{\varphi}_{i}$ (for $i=1,\ldots,N-1$), as depicted in Figure 7. We assume that the reference R-charges of all these chiral multiplets are zero.
@@
$\widetilde{K}_{ab}\equiv\begin{cases}1&\text{if }a=b-1,\quad b=1,\ldots,2N-1,\\[2.0pt] -1&\text{if }a=b-1,\quad b=2N-1,\ldots,4(N-1),\\[2.0pt] -1&\text{if }a=b-N+1,\quad b=1,\ldots,N-1,\\[2.0pt] \dfrac{1}{2}&\text{if }a=b,\quad b=2N-1,\ldots,4(N-1),\\[2.0pt] 0&\text{otherwise}.\end{cases}$ (4.7)
@@
Figure 7: The quiver diagram for the $i$-th set of gauge and matter multiplets included in the CS matter theory for $(A_{2},A_{N-1})$ (for NN coprime to 33). Each circle stands for a $U(1)$ gauge group. Each arrow connecting two circles stands for a bifundamental chiral multiplet, and each arrow from a box to a circle stands for a chiral multiplet charged under the $U(1)$ corresponding to the circle. The subscript of $U(1)$ in each circle stands for the fugacity for the corresponding topological symmetry. This diagram can be read off from (4.4).
@@ 4.3 Monopole superpotential
Table 2: The magnetic charge for monopole operators in the $i$-th sector, $i=1,\dots,N-1$. The monopole operators in the $i$-th sector carry no charge under the gauge groups of the other sectors.
The non-vanishing fugacities (4.9) for the topological symmetries imply that the correct R-symmetry is a linear combination of the reference R-symmetry and the topological symmetries. This suggests that a superpotential involving monopole operators is turned on. From the formula for the gauge charge of monopole operator given in appendix C, we find that the following six types of dressed monopole operators satisfy the gauge-invariance condition $Q_{a}[\mathcal{O}_{(n,m)}]=0$:
@@
Note that, when the above terms are added to the superpotential, the reference R-symmetry is broken since the monopole operators are charged under the R-symmetry (C.2). However, the following linear combination of the reference R-symmetry and topological symmetries is preserved:
@@
$W$ $=\sum_{i=1}^{N-1}\Bigl({\phi}^{(1)}_{i-1}{V}^{(1)}_{i}\phi^{(1)}_{i+1}+{\phi}^{(2)}_{i-1}{V}^{(2)}_{i}\phi^{(2)}_{i+1}+\widetilde{\phi}^{(1)}_{i-1}\widetilde{V}_{i}^{(1)}\widetilde{\phi}^{(1)}_{i+1}$
@@ 4.4 Residual global symmetry
$W$ $\to\sum_{k=1}^{N-1}\left(e^{i\beta_{k}}\varphi_{k-1}V_{k}\,\varphi_{k+1}+e^{i\left(-\frac{\zeta^{(1)}_{k}+\zeta^{(2)}_{k}}{2}+\widetilde{\beta}_{k}\right)}\widetilde{\varphi}_{k-1}\widetilde{V}_{k}\,\widetilde{\varphi}_{k+1}\right)$
$+\sum_{k=1}^{N-1}\left(e^{i\left(-\frac{\zeta_{k}^{(1)}}{2}-\alpha^{(1)}_{k}-\beta_{k}\right)}\phi_{k-1}^{(1)}V_{k}^{(1)}\phi_{k+1}^{(1)}+e^{i\left(-\frac{\zeta^{(2)}_{k}}{2}-\alpha^{(2)}_{k}-\beta_{k}\right)}\phi_{k-1}^{(2)}V_{k}^{(2)}\phi_{k+1}^{(2)}\right)$